OpenStudy (anonymous):

Sean currently has an account balance of $3,981.04. He opened the account five years ago with a deposit of $3,760.19. If the interest compounds daily, what is the interest rate on the account?

OpenStudy (anonymous):

0.6%?

OpenStudy (helder_edwin):

if the interest rate is r% per year then after five years you should have \[ \large 3981.04=3760.19(1+r/100)^5 \]

OpenStudy (helder_edwin):

solve for r

OpenStudy (anonymous):

100.1?

OpenStudy (helder_edwin):

no

OpenStudy (helder_edwin):

solving for \(r\): \[ \large 100\cdot\left(\sqrt[5]{\frac{3981.04}{3760.19}}-1\right)=r \]

OpenStudy (anonymous):

\[100*(5\sqrt1.05 -1)=r\] 100*(5.12-1)=r 100*4.14=r 414=r??!

OpenStudy (helder_edwin):

no

OpenStudy (helder_edwin):

use a calculator. the fraction inside the root is not 1.05

OpenStudy (anonymous):

idk. i give up

OpenStudy (helder_edwin):

r=1.1480

OpenStudy (anonymous):

thanks so much. im not usually this thick...must study more XD